RC meeting and Run Plan, Tuesday Sept. 12, 2023
Return to the RC Meeting Calendar
RC Meeting Information
Time:at 3:30pm today
- Location: Counting house conference room
- ZoomGov Meeting Connection Information
Readiness for Beam
- CG: Plan to lock up the hall by 5PM today; MCC will tune beam into the hall on SWING (after laser work complete - if needed); we will staff OWL and if they are ready we will start the CW beam restoration procedures and checks over OWL. Bill Henry will be on-call to provide support for the CW portion of the checks (target protection checks, ion chamber checks) - shift crew should get an estimate from MCC about timeline and call Bill 30 min ahead of time.
- Outstanding issues:
- laser polarization checks: after MCC sweeps the hall and takes us into controlled access Jack plans to go back in and do polarization checks until 6PM.
- Moller quads: there is a ground leak from the LCW system. Plan is to let them flush (they are NOT needed for optics, but needed for a Moller measurement) and we will check on them during an opportunistic down or when we absolutely need to have a Moller measurement
- Moller dipole: the magnet needs to be trained for ~8h before a measurement can be done. This prevents beam delivery so we need to be careful about scheduling this.
- SBS Left rack: breaker replaced and the setup was looked over by Joe Beaufait. At an opportune time we need facilities to fix the type of breaker they are using for our two phase supply. Alex and Ellen are evening out the power draw among the two phases.
- CG: OWL on Wed only has a SL; we need a TO
- CG: check people taking multiple shifts a day
- We need to sort out walkthroughs better moving fwd
For each subsystem we ask that the responsible person go over the following points:
- The overall status of your subsystem
- Will your subsystem be ready well before Hall lock-up on Friday or at the latest by Monday morning?
- Does your system being ready depend on your team only or are outside resources required?
- SBS magnet and Bigbite magnets ramp and hold
- works; acid flush rescheduled for the long down
- Corrector magnets:
- Target status (magnets, measurements, laser, installation, alignment, polarization, power etc)
- Alignment, laser checks, clean up happening now.
- HCal (Jiwan Poudel)
- Bigbite (Arun Tadepalli)
- AP: some areas of low efficiency in the BigBite GEMs from cosmics run; experts will need to look into it
- Counting House Alarms (Bill Henry)
- Bill tested the alarms and instructions for shift crew response in the wiki (needs further revision?)
- DAQ (Alex Camsonne)
- one channel of GRINCH has an issue => fixed by Eric and Todd; PMT base was replaced
- SBS GEMs are hanging the DAQ (Vimukthi) => should be ok now;
- GEMs (Holly Szumila-Vance)
- ChG: Ellen has a mass flow controller and needs to find some time to install it (not a show stopper - gas mixture changes from 79%Ar to 70%) => Chandan can also change the MFC; this is outside of the hall can be done later
- Moller polarimeter (Don Jones)
- DJ: training started; will be interrupted 1h before beam delivery
- waiting on replay
Preliminary run plan
- First, bring the beam safely into the hall (Beam recovery procedure - Jiwan and William Henry https://sbs.jlab.org/wiki/index.php/HOW_TOs#Beam_Recovery_Procedure)
Q: At what stage is the SBS turned ON? WH: will post links to the corrector checkout
- BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)
- Target performance assessment by target group - Arun (Why hasn’t the target reached xx% polarization?), tell us when we will get there (Target group - Arun) and plan for those hours in between
- Take production data!
List of items in case target isn’t ready:
- Take data on reference cell at 10 atm H2 (portions that add up to 1 day)
- BigBite Sieve slit
- with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
- 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
- Moller measurement depending on availability of people and results from Hall C
In person: Ciprian G, Alexandre C, Chandan G, Hanjie L, Zak R, Don J, Eric F, Jimmy C
Online: Todd A, Ole H, Vimukthi G, Jiwan P, Sanghwa P, Sean J, Silviu CD, Andrew P, Bryan M,